OpenStudy (anonymous):

Change to a decimal: 14 1/2 %. 0.145 0.0145 14.5

OpenStudy (anonymous):

14/100 = .145 just move the decimal over two places

OpenStudy (anonymous):

1/5 = .5

OpenStudy (anonymous):

1/2*****

OpenStudy (anonymous):

wah ?

OpenStudy (anonymous):

its 14 1/2 %

OpenStudy (anonymous):

= (14 1/2%)/100% = (29/2)/100 = 29/200 or 0.145 Answer: A. 0.145

OpenStudy (anonymous):

so the answer is A. 0.145 ☻

OpenStudy (anonymous):

What?the answer is 0.145? wierd

OpenStudy (anonymous):

What?the answer is 0.145? wierd

OpenStudy (anonymous):

What?the answer is 0.145? wierd

OpenStudy (anonymous):

What?the answer is 0.145? wierd

OpenStudy (anonymous):

What?the answer is 0.145? wierd